Suppose two inputs are perfect substitutes. In order to minimize production costs, the producer will simply have to employ the cheapest input.

a. True b. False
Answer #1

Answer

Flase

The producer hires the input which provides the higher per dollar marginal product

Ex.

MPL=10 and w=10, MPK=5 and r=7.5

per dollar MP=MP/price of input

per dollar MPL=10/10=1

per dollar MPK=5/7.5=0.666666667=0.67

the capital is cheaper but the produce hires labor because the labor provides higher per dollar marginal product.